# Security Approaches for Data Provenance in the Internet of Things: A Systematic Literature Review

### v1

@@ Abstract.
The Internet of Things (IoT) relies on resource-constrained devices deployed in unprotected environments. Different types of risks may be faced during data transmission in single-hop and multi-hop scenarios. Addressing these vulnerabilities is crucial. A systematic literature review of data provenance in IoT is presented, exploring existing techniques, practical implementations, security requirements, and performance metrics. Respective contributions and shortcomings are compared. A taxonomy related to the development of data provenance in IoT is proposed. Open issues are identified, and future research directions are presented, providing useful insights for the evolution of data provenance research in the context of the IoT.
Internet of Things, Data Provenance, Intrusion Detection, Data Integrity, Cyber Security, Trustworthiness.
ccs: General and reference Surveys and overviewsccs: Security and privacy Trustworthinessccs: Security and privacy IoT securityccs: Security and privacy Secure provenance
Acknowledgments
The authors acknowledge the funding obtained from the PID2021-125962OB-C31 “SECURING” project funded by the Ministerio de Ciencia e Innovación, la Agencia Estatal de Investigación and the European Regional Development Fund (ERDF), as well as the ARTEMISA International Chair of Cybersecurity and the DANGER Strategic Project of Cybersecurity, both funded by the Spanish National Institute of Cybersecurity through the European Union – NextGenerationEU and the Recovery, Transformation and Resilience Plan. The authors also acknowledge the support from IMT’s CyberCNI chair (Cybersecurity for Critical Networked Infrastructures, cf. https://cybercni.fr/).
@@ 1. Introduction
Figure 1. Workflow of this paper.
An intelligent system known as the Internet of Things (IoT) comprises a network infrastructure of physical objects that are interconnected, enabling the exchange and collection of data from diverse sources and destinations. The objects are designed to transfer data captured from the environment over an insecure internet connection, where it is processed, managed, and analyzed through various technologies (Ammar et al., 2018; Botta et al., 2016). IoT increases the level of network vulnerabilities as it becomes more accessible to attackers. Despite this, the intelligent system is widely applied in different fields such as industrial control, healthcare systems, home appliances, car automation, and environmental monitoring (Lazarescu, 2017; Manikandan and Subha, 2018). These systems rely on various sensors for data acquisition and transmission, which are resource constrained due to deployment in an unprotected physical environment. IoT networks operate via single-hop and multi-hop structures, with data transmission vulnerable to tampering (Gubbi et al., 2013). Furthermore, the data generated from a huge number of sensor node sources is subjected to in network processing by intermediate nodes during transmission to a base station, and this processed data is used in the decision making processes, making it critical to evaluate data trustworthiness, detect attacks during data generation and processing, and identify nodes and actions responsible for those attacks (Sultana et al., 2015). IoT systems lack the capacity to perform identity management, ensure data trustworthiness, detect abnormal behavior, and control unauthorized data access, becoming vulnerable to cyber attacks (Hu et al., 2020). Data provenance solutions have been proposed aiming to contribute to overcome the aforementioned security issues in IoT systems. Initially, data provenance was used for heterogeneous database systems (Wang and Madnick, 1990). It was originally introduced to document the origin, history, derivation, creation and usage of some entity, which is also known as lineage (Wang et al., 2016b). It explains how data evolves from process to another. Many researchers used provenance in the field of information technology, which is defined as the process of recording the history of origin, evolution, process activities and manipulation of data over time (Buneman et al., 2001a; Simmhan et al., 2005; Zafar et al., 2017). Researchers introduce data provenance as a solution to provide a number of required features to ensure trustworthiness, integrity, data quality, confidentiality, availability and other security requirements. In this regard, data provenance is presented in many domains such as file systems, databases, cloud computing, IoT, distributed networks and many more domains. This research focuses on data provenance within the IoT and Wireless Sensor Networks (WSN) domains.
@@ 1.2   Major Contributions — The main contributions of our survey are summarized below:
1.3   Paper Organization — Section 2 presents the review methodology used to conduct our survey. Section 3 provides an overview on IoT network and data provenance. Section 4 presents existing surveys on data provenance in IoT. Section 5 introduces existing data provenance approaches in IoT networks, with a comparative analysis. Section 6 provides a discussion. Section 7 presents open issues and research challenges. Section 8 concludes the paper.
@@ 2. Review methodology
RQ5 What are the advantages and limitations of the proposed techniques for data provenance in IoT in the studied literature? The answer will help specify the contributions and shortcomings of the studied state of the art Section 5 and  6
@@
2.2   Search strategy — To address the research questions, a search process was conducted across six electronic publication databases. The selected databases are shown in Table 2. These databases are widely recognized and widely used in the fields of computer science and engineering, offering comprehensive coverage of the relevant literature. The automated search process aimed to retrieve relevant studies from these databases (Chen et al., 2010; Malheiros et al., 2007). The set of keywords used to find these studies include the following: Internet of Things, Wireless Sensor Network, Data Provenance, and Secure Provenance. The search strategy in the mentioned electronic databases was performed using the following query (and considering synonyms and alternative terms [^1] The primary keywords were linked together using the logical operator “AND” to establish the search parameters. Similarly, the variations and synonyms were connected using the logical operator “OR” to broaden the scope of the search. ):
@@
2.3   Selection criteria — To assess the relevance of each primary study in relation to the defined research questions, different selection criteria were applied. The objective was to include studies that have the potential to address the research questions and exclude those that do not contribute to answering them. The criteria is used as a means of determining the suitability of each study for inclusion in the analysis. We have considered several inclusion and exclusion criteria as follows:
@@
EC2: The article is not focused on data provenance in the domain of IoT or WSN.
@@ 3. Data provenance in IoT networks
3.1   Overview of IoT networks — The realization of the IoT concept in practical terms can be achieved by combining various enabling technologies. In this survey, we direct our attention towards sensor networks, which stand out as a highly used technology across a wide range of applications. Sensor networks have been proposed for numerous scenarios, including environmental monitoring, e-health, intelligent transportation systems, military applications, and industrial monitoring , among others. These networks consist of a number of sensing nodes that communicate using a wireless multi-hop model. Typically, nodes transmit captured data to a limited number of intermediate forwarding nodes (Atzori et al., 2010). In recent years, scientists have conducted extensive research on sensor networks, examining various challenges across different layers of the network protocol stack (Akyildiz et al., 2002). It faces significant vulnerability to attacks due to several factors. Firstly, many of its components are often in stand-by mode for extended periods, making them easy targets for physical attacks. Secondly, the widespread use of wireless communication in the IoT makes eavesdropping a simple task. Lastly, most IoT components possess limited capabilities in terms of energy and computing resources, which prevents the implementation of complex security schemes. Specifically, trustworthiness and data integrity present major security concerns (Atzori et al., 2010). The presence of diverse data sources requires the establishment of trustworthiness for the data, ensuring that only reliable information is taken into account during the decision-making process. In addition, data integrity solutions play an important role in ensuring that any unauthorized modifications to data being transmitted by source nodes or intermediate nodes are detected by the system, preventing adversaries from tampering with the information. For these concerns, new solutions were proposed to ensure the trustworthiness of data, such that any decision process is based on trustworthy data. Data provenance is introduced as one of these effective solution that can be used in constrained networks to ensure trustworthiness of sensed data. Data provenance techniques must consider additional methods to tackle the problem of provenance integrity and data packet integrity.
Figure 3. Taxonomy of Data Provenance in IoT networks.
@@
3.2   What is data provenance? — The definition of provenance has undergone changes over time, adapting to different application contexts. In database systems, provenance was traditionally referred to as lineage and focused on identifying the source of data resulting from query processing, commonly referred to as data provenance (Pan et al., 2023; Moreau et al., 2018). Data lineage was initially formalized by Cui et al. (2000), specifically in the field of relational databases. It aimed to trace each tuple "$t"$ in the input tables that played a role in generating the output of a query from the database (Cheney et al., 2009). Provenance, also referred to as pedigree, or genealogy, is a form of metadata that documents the origin and utilization of a given entity (Simmhan et al., 2005; Glavic and Dittrich, 2007; Buneman et al., 2001b). In the field of information technology, provenance considers data as the counterpart or reflection of an art object. In this field, researchers also define provenance as the complete information about the entire process of data generation and evolution over time. This includes capturing the static origins of data as well as tracking their dynamic changes throughout their transmission process (Gao et al., 2010). As a summary, provenance holds the capability to provide insights into the what, where, when, how, and why aspects of a given data object.
@@
3.3   Data provenance integration with IoT — The interconnectivity of the IoT has brought significant improvements to a wide range of different application scenarios. It has also introduced security and privacy concerns due to data transmission over open networks. Attackers can easily inject malicious data into the transmission path and maliciously manipulate data without any notice from the entities of the networks, thereby compromising the integrity and trustworthiness of the data. Additionally, the rapid spread of malicious data can cause catastrophic failures. Moreover, the sharing, transmission, and processing of data leads to the risk of user privacy breaches. Data provenance, which records the origin and processing history of data, presents a potential solution to address these aforementioned issues. Provenance information can record the original source node which produced or forwarded data packet and the operations which the data went through. In many cases, provenance can provide the time and location of the entities that engaged in any operational procedure on data. We provide researchers with a proposed taxonomy for the different aspects of data provenance in IoT as shown in Figure 3.
@@
  • $\mathrm{HOST}(\cdot)$: The function that assigns a host node to each network node in $V$.$\mathrm{HOST}(v_{i})$ denotes the host node of the network node $v_{i}$.
@@
3.4   Data provenance storage —  Efficiently managing the increasing granularity of captured information in provenance, as the number of sensor nodes in a network grows, is a crucial demand. In IoT, the provenance expands rapidly due to the increase participation of forwarding nodes. The elements of provenance information vary across applications and methods, depending on the specific requirements fulfilled by the proposed solution. Furthermore, gathering detailed information about data generation, processing, and forwarding enables the system to ensure diverse security requirements. Therefore, it is essential to store this data efficiently to minimize bandwidth, storage, and energy overhead. Provenance information can be stored using four main storage techniques: local database, blockchain, in-packet storage, and cloud-based.
@@
    Local Database: In the local database system, provenance information is stored either in a distributed or centralized database, depending on the deployed solution’s application. This storage system has challenges regarding storing and querying flexibility, as well as how IoT nodes store data information at each forwarding node. Figure 5 shows an example of the use of a local database in an IoT network.
@@
    Blockchain-based: Blockchain, a peer-to-peer network’s decentralized ledger maintained by all peers, which offers distributed data storage and has been applied to establish data provenance by recording data processes as blockchain transactions (Salman et al., 2019). The large number of data packets generated from different source nodes in the IoT network complicates the use of blockchain for storing provenance records, as each record becomes a transaction. This complexity introduces challenges and limitations when applying blockchain for provenance storage. An example of connecting a blockchain to an IoT network for data provenance is shown in Figure 5.
@@
3.4.1   Data provenance categories. The main objective when implementing a new data provenance method is to securely store and transmit provenance information. To do so, many challenges in term of security and privacy need to be addressed, which is related to provenance manipulation in collection, storing, and transmission. It is also essential for the designed system to consider the security requirements to overcome such challenging issues (Liu and Wu, 2020; Pan et al., 2023). In this context, secure provenance solutions can be divided into the following categories: Cryptography-based, Digital Watermarking, Bloom filters, Physical Unclonable Functions, Fingerprints, Blockchain-based, Frameworks with storing methods, Data Santization, Lexical chaining, Graph-based, Path difference and Logging-based. These categories includes different security techniques that may be combined in some applications to assure data provenance in different IoT scenarios. These methods rely on many factors to be developed such as network resources, IoT application, needed security requirements, provenance storage approach, energy utilization, storage overhead, attack types, and network architecture. The categorization of security techniques for data provenance is shown in the proposed taxonomy in Figure 3. Moreover, a detailed definition and overview on the different categories with their respective techniques is provided in Section 5.
3.5   Security Requirements — Introducing the challenges of ensuring data trustworthiness in such limited and constrained networks have set a number of security requirements that should be satisfied to have a robust system against different types of attacks. These requirements differ from one scenario to another and are based on the application of the IoT system. The requirements are described as follows:
  • Privacy: Provide extra guarantees that the sources of this provenance data are not tracked by unauthorized parties. In many cases, the provenance information is more important than the data itself. It requires protecting privacy of such data.

### v2

@@ Abstract.
The Internet of Things (IoT) relies on resource-constrained devices deployed in unprotected environments. Given their constrained nature, IoT systems are vulnerable to security attacks. Data provenance, which tracks the origin and flow of data, provides a potential solution to guarantee data security, including trustworthiness, confidentiality, integrity, and availability in IoT systems. Different types of risks may be faced during data transmission in single-hop and multi-hop scenarios, particularly due to the interconnectivity of IoT systems, which introduces security and privacy concerns. Attackers can inject malicious data or manipulate data without notice, compromising data integrity and trustworthiness. Data provenance offers a way to record the origin, history, and handling of data to address these vulnerabilities. A systematic literature review of data provenance in IoT is presented, exploring existing techniques, practical implementations, security requirements, and performance metrics. Respective contributions and shortcomings are compared. A taxonomy related to the development of data provenance in IoT is proposed. Open issues are identified, and future research directions are presented, providing useful insights for the evolution of data provenance research in the context of the IoT.
Internet of Things, Data Provenance, Intrusion Detection, Data Integrity, Cyber Security
copyright: acmcopyrightbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: General and reference Surveys and overviewsccs: Security and privacy Trustworthinessccs: Security and privacy IoT securityccs: Security and privacy Secure provenance
Acknowledgments: Authors acknowledge funding from the PID2021-125962OB-C31 “SECURING” project funded by the Ministerio de Ciencia e Innovación, la Agencia Estatal de Investigación and the European Regional Development Fund (ERDF), as well as the ARTEMISA International Chair of Cybersecurity and the DANGER Strategic Project of Cybersecurity, both funded by the Spanish National Institute of Cybersecurity through the European Union – NextGenerationEU and the Recovery, Transformation and Resilience Plan; and IMT’s CyberCNI chair (Cybersecurity for Critical Networked Infrastructures, cf. https://cybercni.fr/).
@@ 1. Introduction
The Internet of Things (IoT) comprises a network of interconnected physical objects, enabling the exchange and collection of data from diverse sources. This infrastructure facilitates data transfer from various environments over an insecure internet connection, where it is processed, managed, and analyzed through various technologies (Ammar et al., 2018; Botta et al., 2016). However, IoT increases network vulnerabilities, as it often lacks mechanisms for identity management, data trustworthiness, and access control, making it accessible to attackers. IoT applications span across industrial control, healthcare, home automation, car systems, and environmental monitoring (Lazarescu, 2017; Manikandan and Subha, 2018; Gubbi et al., 2013). Moreover, data generated by numerous sensor nodes is subjected to in-network processing during transmission to a base station, where provenance, or the tracking of data origin and evolution, becomes critical to ensure data trustworthiness in decision-making processes (Sultana et al., 2015).
To address IoT security challenges, data provenance is introduced as a solution to trace and verify the history of data transactions. Initially developed for heterogeneous database systems (Wang and Madnick, 1990), data provenance documents the origin, evolution, and transformations applied to data entities, also known as lineage (Wang et al., 2016b). In IoT and other domains, data provenance is defined as the process of recording the history of data origin, evolution, and activities, providing essential security features such as trustworthiness, integrity, and quality assurance (Buneman et al., 2001a; Simmhan et al., 2005; Zafar et al., 2017). The objective of data provenance is not only to ensure data quality but also to address specific security requirements, including confidentiality, availability, and the prevention of unauthorized access. This paper proposes a taxonomy aimed at providing an in-depth understanding of data provenance in IoT by classifying data provenance solutions across three primary aspects: encoding methods, storage mechanisms, and security attack resilience. Provenance encoding refers to techniques used to encode provenance information into IoT systems, enabling efficient tracking of data origins. Provenance storage mechanisms focus on various methods to store and retrieve provenance data, while security attacks represent threats against both data and provenance information in IoT networks, addressing mechanisms to ensure data integrity and authenticity in adversarial environments. With provenance applied in domains like file systems, databases, cloud computing, and IoT networks, this research specifically focuses on its application within the IoT and Wireless Sensor Networks (WSN) domains, emphasizing how provenance can be leveraged to enhance data security and authenticity in these environments by providing researchers with a deep understanding of data provenance and offering insights into research challenges, open issues, and future directions.
@@ 1.2   Major Contributions — The main contributions of our survey are summarized below:
Figure 1. Workflow of this paper.
1.3   Paper Organization — Section 2 presents the review methodology used to conduct our survey. Section 3 provides an overview on IoT network and data provenance. Section 4 provides the mechanisms for data provenance storage systems. The security attacks on IoT networks is introduced in Section 5. Section 6 presents existing surveys on data provenance in IoT. Section 7 introduces existing data provenance approaches in IoT networks, with a comparative analysis. Section 8 provides a discussion. Section 9 presents open issues and research challenges. Section 10 concludes the paper.
@@ 2. Review methodology
RQ5 What are the advantages and limitations of the proposed techniques for data provenance in IoT in the studied literature? The answer will help specify the contributions and shortcomings of the studied state of the art Section 7 and  8
@@
2.2   Search strategy — To address the research questions, a search process was conducted across six electronic publication databases. The selected databases are shown in Table 2. These databases are widely recognized and widely used in the fields of computer science and engineering, offering extensive coverage of the relevant literature. The automated search process aimed to retrieve relevant studies from these databases (Chen et al., 2010; Malheiros et al., 2007). The set of keywords used to find these studies include the following: Internet of Things, Wireless Sensor Network, Data Provenance, and Secure Provenance. The search strategy in the mentioned electronic databases was performed using the following query (and considering synonyms and alternative terms [^1] The primary keywords were linked together using the logical operator “AND” to establish the search parameters. Similarly, the variations and synonyms were connected using the logical operator “OR” to broaden the scope of the search. ):
@@
2.3   Selection criteria — To assess the relevance of each primary study in relation to the defined research questions, different selection criteria were applied. The objective was to include studies that have the potential to address the research questions and exclude those that do not contribute to answering them. The criteria is used as a means of determining the suitability of each study for inclusion in the analysis. We have considered several inclusion and exclusion criteria show in Table 3.
Table 3. Inclusion and exclusion criteria used for inclusion in our analysis.
@@
EC2: The article is not focused on data provenance in the domain of IoT or Wireless Sensor Networks (WSN).
@@ 3. Data provenance in IoT networks
3.1   Overview of IoT networks — The realization of the IoT concept in practical terms can be achieved by combining various enabling technologies. In this survey, we direct our attention towards sensor networks, which stand out as a highly used technology across a wide range of applications. Sensor networks have been proposed for numerous scenarios, including environmental monitoring, e-health, transportation systems, military applications, and industrial monitoring , among others. These networks consist of a number of sensing nodes that communicate using a wireless multi-hop model. Typically, nodes transmit captured data to a limited number of intermediate forwarding nodes (Atzori et al., 2010). In recent years, scientists have conducted extensive research on sensor networks, examining various challenges across different layers of the network protocol stack (Akyildiz et al., 2002). It faces significant vulnerability to attacks due to several factors. Firstly, many of its components are often in stand-by mode for extended periods, making them easy targets for physical attacks. Secondly, the widespread use of wireless communication in the IoT makes eavesdropping a simple task. Lastly, most IoT components possess limited capabilities in terms of energy and computing resources, which prevents the implementation of complex security schemes. Specifically, trustworthiness and data integrity present major security concerns (Atzori et al., 2010). The presence of diverse data sources requires the establishment of trustworthiness for the data, ensuring that only reliable information is taken into account during the decision-making process. In addition, data integrity solutions play an important role in ensuring that any unauthorized modifications to data being transmitted by source nodes or intermediate nodes are detected by the system, preventing adversaries from tampering with the information. For these concerns, new solutions were proposed to ensure the trustworthiness of data, such that any decision process is based on trustworthy data. Data provenance is introduced as one of these effective solution that can be used in constrained networks to ensure trustworthiness of sensed data. Data provenance techniques must consider additional methods to tackle the problem of provenance integrity and data packet integrity.
Figure 3. Proposed Taxonomy of Data Provenance in IoT networks.
@@
3.2   What is data provenance? — The definition of provenance has undergone changes over time, adapting to different application contexts. In database systems, provenance was traditionally referred to as lineage and focused on identifying the source of data resulting from query processing, commonly referred to as data provenance (Pan et al., 2023; Moreau et al., 2018). Data lineage was initially formalized by Cui et al. (2000), specifically in the field of relational databases. It aimed to trace each tuple "$t"$ in the input tables that played a role in generating the output of a query from the database (Cheney et al., 2009). Provenance, also referred to as pedigree, or genealogy, is a form of metadata that documents the origin and use of a given entity (Simmhan et al., 2005; Glavic and Dittrich, 2007; Buneman et al., 2001b). In the field of information technology, provenance considers data as the counterpart or reflection of an art object. In this field, researchers also define provenance as the complete information about the entire process of data generation and evolution over time. This includes capturing the static origins of data as well as tracking their dynamic changes throughout their transmission process (Gao et al., 2010). As a summary, provenance holds the capability to provide insights into the what, where, when, how, and why aspects of a given data object.
@@
3.3   Data provenance integration with IoT — The interconnectivity of the IoT has brought significant improvements to a wide range of different application scenarios. It has also introduced security and privacy concerns due to data transmission over open networks. Attackers can easily inject malicious data into the transmission path and maliciously manipulate data without any notice from the entities of the networks, thereby compromising the integrity and trustworthiness of the data. For example, in 2016, the Mirai botnet attack exploited unsecured IoT devices, resulting in widespread disruptions of major websites, demonstrating the potential for privacy breaches through compromised devices (Pankov, 2023). Additionally, the rapid spread of malicious data can cause catastrophic failures. Moreover, the sharing, transmission, and processing of data leads to the risk of user privacy breaches. Unauthorized access to smart home systems can allow intruders to activate webcams and voice-controlled devices or manipulate household appliances like ovens and electric stoves, thereby compromising the security of the entire smart home ecosystem. Such breaches not only jeopardize the functionality of these devices but also pose significant risks to user privacy, as unauthorized individuals may gain access to sensitive personal information or monitor household activities without consent (Buil-Gil et al., 2023). Data provenance, which records the origin and processing history of data, presents a potential solution to address these aforementioned issues. Provenance information can record the original source node which produced or forwarded data packet and the operations which the data went through. In many cases, provenance can provide the time and location of the entities that engaged in any operational procedure on data. We provide researchers with a proposed taxonomy for the different aspects of data provenance in IoT as shown in Figure 3.
@@
  • $\mathrm{HOST}(\cdot)$ assigns a host node to each network node in $V$, where $\mathrm{HOST}(v_{i})$ denotes the host node of the network node $v_{i}$.
@@
3.4   Data provenance categories – The main objective when implementing a new data provenance method is to securely store and transmit provenance information. To do so, many challenges in term of security and privacy need to be addressed, which is related to provenance manipulation in collection, storing, and transmission. It is also essential for the designed system to consider the security requirements to overcome such challenging issues (Liu and Wu, 2020; Pan et al., 2023). In this context, secure provenance solutions can be divided into the following categories: Cryptography-based, Digital Watermarking, Bloom filters, Physical Unclonable Functions, Fingerprints, Blockchain-based, Frameworks with storing methods, Data Santization, Lexical chaining, Graph-based, Path difference and Logging-based. These categories includes different security techniques that may be combined in some applications to assure data provenance in different IoT scenarios. These methods rely on many factors to be developed such as network resources, IoT application, needed security requirements, provenance storage approach, energy usage, storage overhead, attack types, and network architecture. The categorization of security techniques for data provenance is shown in the proposed taxonomy in Figure 3. Moreover, a detailed definition and overview on the different categories with their respective techniques is provided in Section 7.
3.5   Security Requirements — Introducing the challenges of ensuring data trustworthiness in such limited and constrained networks have set a number of security requirements that should be satisfied to have a robust system against different types of attacks. The security requirements were identified through an in-depth analysis of related work (detailed in Table 15 of Appendix A) and by considering the specific characteristics and constraints of IoT environments, such as limited computational resources and energy efficiency. We also evaluated common attack vectors and vulnerabilities in IoT systems to ensure that the requirements address both theoretical and practical security concerns. We can observe that none of the selected papers in the systematic literature review satisfy all of the security requirements necessary for a fully robust provenance system as shown in Table 7. For example, some papers focus on data integrity and confidentiality but lack provisions for ensuring freshness and unforgeability. Others address availability but fail to account for non-repudiation or privacy. This gap highlights the need for a robust solution that encompasses all the key security aspects to effectively protect provenance information in IoT networks. These requirements differ from one scenario to another and are based on the application of the IoT system. The requirements are described as follows:
  • Privacy: It provides extra guarantees that the sources of this provenance data are not tracked by unauthorized parties. In many cases, the provenance information is more important than the data itself. It requires protecting privacy of such data.
4. Data provenance storage
Efficiently managing the increasing granularity of captured information in provenance, as the number of sensor nodes in a network grows, is a crucial demand. In IoT, the provenance expands rapidly due to the increase participation of forwarding nodes. The elements of provenance information vary across applications and methods, depending on the specific requirements fulfilled by the proposed solution. Gathering detailed information about data generation, processing, and forwarding enables the system to ensure diverse security requirements. Thus, it is essential to store this data efficiently to minimize bandwidth, storage, and energy overhead. Provenance information can be stored using four main storage techniques: local database, blockchain, in-packet storage, and cloud-based.
@@ 4. Data provenance storage
    Local Database: In the local database system, provenance information is stored either in a distributed or centralized database, depending on the deployed solution’s application. However, local storage is typically located at the edge in IoT networks for provenance information storage, such as in an edge server, and not directly at resource-constrained IoT nodes. This distinction is important, as the storage at the edge does not require an internet connection but is still more capable than individual IoT devices. Moreover, as the provenance record grows with the number of forwarding hops, it is not feasible to store it in constrained IoT nodes. By storing the information at the edge, it is possible to accommodate larger provenance data without impacting the performance of the IoT devices. This approach also enhances scalability, as more resources are available at the edge to handle increasing data volumes (Zafar et al., 2017). This storage system has challenges regarding storing and querying flexibility, as well as how IoT nodes store data information at each forwarding node. Figure 5 shows an example of the use of a local database in an IoT network.
@@
    Blockchain-based: Blockchain is a peer-to-peer network’s decentralized ledger maintained by all peers, which offers distributed data storage and has been applied to establish data provenance by recording data processes as blockchain transactions (Salman et al., 2019). The large number of data packets generated from different source nodes in the IoT network complicates the use of blockchain for storing provenance records, as each record becomes a transaction. This complexity introduces challenges and limitations when applying blockchain for provenance storage. However, in addition to traditional internet-based blockchain solutions, edge-deployed blockchain frameworks, such as Hyperledger Fabric, can also be used to store provenance information closer to the data sources. An example of connecting a blockchain to an IoT network for data provenance is shown in Figure 5.
@@
5. Security Attacks